### v1

# Efficient parallel scheduling for sparse triangular solvers

Abstract.
We develop and analyze new scheduling algorithms for solving sparse triangular linear systems (SpTRSV) in parallel. Our approach, which we call barrier list scheduling, produces highly efficient synchronous schedules for the forward- and backward-substitution algorithm. Compared to state-of-the-art baselines HDagg [ZCL+22] and SpMP [PSSD14], we achieve a 3.24×3.24$\times$3.24 × and 1.45×1.45$\times$1.45 × geometric-mean speed-up, respectively. We achieve this by obtaining an up to $\times1$ geometric-mean reduction in the number of synchronization barriers over HDagg, whilst maintaining a balanced workload, and by applying a matrix reordering step for locality. We show that our improvements are consistent across a variety of input matrices and hardware architectures.
Key words and phrases:
Sparse triangular linear system solve, SpTRSV, SpTRSM, forward- and backward-substitution algorithm, barrier list scheduler, synchronous parallel algorithm.
22footnotetext: Authors are listed in alphabetical order.
1. Introduction
Systems of linear equations are ubiquitous and solving them fast numerically with high accuracy is essential to engineering, big data analytics, artificial intelligence, and various scientific subjects. Key techniques in scaling to ever larger linear systems have been exploiting the sparsity of non-zero coefficients in modern algorithms, as well as leveraging the multi-core or multi-processor architectures of high-performance computing systems. However, whilst sparsity reduces computational load, the typically irregular distribution of non-zero elements complicates the development of efficient parallel algorithms, as the lack of structure hinders workload balancing and limits the ability to minimize communication between processors.
In this paper, we concern ourselves with solving sparse triangular systems of linear equations (SpTRSV) using parallel machines; i.e., solving a linear system $Lx=b$, where $L$ is a sparse triangular matrix and $b$ is a dense vector. Although solving sparse triangular linear systems mark a special case, it often arises as an important step in procedures solving more general linear systems. Some concrete examples are (sparse) LU, QR, and Cholesky decompositions, Gauss–Seidel, and so forth. Efficient parallel-computation schedules for SpTRSV are of particular importance in applications where the same sparsity pattern is used repeatedly. Such is the case in simulations of various physical systems, for instance, ones that are based on the finite element method on a fixed mesh.
One of the main methods of solving SpTRSV is the forward-/backward-substitution algorithm. An execution of the algorithm on an instance may be captured by a directed acyclic graph (DAG), with the vertices corresponding to the rows of the matrix and directed edges representing dependencies imposed by the non-zero entries, see Figure 1.1. Finding a parallel execution of the forward-/backward-substitution algorithm directly corresponds to solving the parallel-scheduling problem on the corresponding DAG.
abcdefabcdef
(a) Sparse lower triangular matrix ($6\times 6$)
abcdfe
(b) Corresponding DAG
Figure 1.1. A sparse lower triangular matrix (a) and its corresponding DAG for the forward-substitution algorithm (b). Each row of the matrix corresponds to a vertex in the DAG. An edge from vertex $u$ to vertex $v$ exists if and only if there is a non-zero entry in column $u$ of row $v$ in the matrix. The dotted lines in Figure (b) separate the wavefronts of the DAG.
In order to generate an efficient parallel schedule for the algorithm, one needs to:
  1. (i)
    balance workload across machines, and
  2. (ii)
    limit coordination overhead.
Satisfying both of these needs simultaneously has proven to be challenging due to the irregular interdependence of computed values and the fine-grained nature of the problem. Early algorithms include so-called wavefront schedulers [AS89, Sal90], which repeatedly schedule all computations whose prerequisites are met, known as the wavefronts, cf. Figure 1.1(b), followed by a synchronization barrier. They, however, suffer from large overhead stemming from frequent global synchronization [PSSD14]. Similarly, early asynchronous approaches such as self-scheduling [SMB88] had the drawback of incurring overheads due to numerous fine-grained synchronizations [RG92].
In a breakthrough paper, Park et al. [PSSD14] reduced coordination overhead by combining these earlier ideas. Their scheduler SpMP, which remains a competitive baseline to date, is in essence an asynchronous wavefront scheduler: it allows machines to move onto the next wavefront if and only if all requisites have already been met for its portion of the next wavefront. They also developed a fast approximate transitive reduction to reduce the number of synchronization points further. An alternate reduction in synchronizations has been made by Yilmaz et al. [YSAU20] by enforcing a bound by which machines may be out of sync.
For synchronous schedulers, efforts have been directed towards increasing the computational load between synchronization barriers, thus decreasing the number of global synchronizations. For instance, Cheshmi et al. [CKSD18] devise such methods for triangular matrices of a special structure, arising in Cholesky decompositions. For general sparse triangular matrices, a state-of-the-art baseline is the recent scheduler HDagg of Zarebavani et al. [ZCL+22]. This algorithm develops efficient schedules by gluing together consecutive wavefronts if and only if a balanced workload can still be maintained and by pre-applying a DAG coarsening technique.
1.1. Our contribution
Our work continues along the same path of reducing the number of synchronization barriers. We put forth barrier list schedulers as an effective synchronous scheduling algorithm archetype for solving sparse triangular linear systems. In particular, we present and analyze two specific algorithms that are both tailored towards the SpTRSV application. In our experiments, we establish that these algorithms produce significantly superior parallel schedules compared to the baseline methods. Specifically, our Funnel Locking algorithm achieves a reduction in execution time of 1.45×1.45$\times$1.45 × compared to SpMP and of 3.24×3.24$\times$3.24 × compared to HDagg, on the SuiteSparse Matrix Collection benchmark [DH11] used by previous studies, see Figure 1.2. On uniformly random matrices, our other algorithm, called Funnel $p$-ivotal path, achieves a similar speed-up of 1.62×1.62$\times$1.62 × compared to SpMP and 1.87×1.87$\times$1.87 × compared to HDagg in execution time.
Figure 1.2. Geometric mean and interquartile ranges of speed-ups over Serial of several algorithms on the SuiteSparse Matrix Collection [DH11] on an Intel x86 machine using 22 cores.
Our algorithms obtain these speed-ups by significantly reducing the number of synchronization barriers required: we report an up to $\times1$ reduction in the number of barriers relative to HDagg, whilst maintaining a good workload balance. The results also show that our schedulers provide consistent improvements over several different computing architectures and types of input matrices. The scheduling algorithms themselves also have a just above linear time complexity, making them viable tools for various applications.
In summary, the main contributions of our paper are:
  • two novel algorithms for generating efficient parallel schedules for SpTRSV execution,
  • adaptations and extensions of previous coarsening and reordering techniques that enhance the schedules,
  • a short theoretical proof showing that our DAG coarsening techniques preserve acyclicity, based on the new concept of cascades, and
  • experiments confirming that the above schedulers achieve significant speed-ups over the SpMP and HDagg baselines, on various architectures and data sets, including orthogonal testing of the individual techniques proposed.
1.1.1. Barrier list schedulers
There is an extensive amount of prior work on parallel DAG scheduling in the literature. A recent survey classifies the most relevant algorithms into two major groups, list schedulers and cluster schedulers [WS18]. List schedulers [Gra69, ACD74, HCAL89, RVG02, MSQ03] typically assign priorities to each vertex and schedule them in a topological order according to these. Such algorithms are often more efficient when the number of cores is limited [WS18].
The idea of list schedulers has also been adapted recently to a setting with barrier synchronization [PAKY24]. We refer to this new approach as barrier list scheduling. Similarly to list scheduling, these algorithms maintain a set of vertices that are ready to be executed, i.e., all their parents have been computed. Whenever a core $p$ becomes free, they assign a ready vertex to $p$ that does not require new communication between cores, with a preference for vertices that can only be executed on the core $p$. The schedulers only introduce a new synchronization barrier when a given fraction of cores become idle.
In our paper, we present two barrier-list-scheduling algorithms for finding efficient parallel schedules for the SpTRSV kernel. We extend the prototype of the barrier-list-scheduling idea by Papp et al. [PAKY24] with various improvements: we refine the rule for inserting a synchronization barrier, we decrease the amount of compute resources wasted before the synchronization barriers, and we apply more advanced data structures for an efficient implementation. More importantly, the key to the behavior of a barrier-list-scheduling algorithm is the priority function, which selects the next vertex to be assigned to a free core whenever there are multiple options. In our two schedulers, these priority functions are tailored specifically towards the SpTRSV application:
  • the $p$-ivotal path scheduler aims to prioritize vertices with long outgoing paths and/or a high-number of successors in the DAG, whilst
  • the Locking scheduler aims to ensure that we can compute the highest possible number of vertices before having to insert a new synchronization barrier.
While both algorithms are similar, our experiments show that Locking performs better on the SuiteSparse Matrix Collection and $p$-ivotal path is superior on the (unstructured) random data sets.
1.1.2. Coarsening
We combine our scheduling algorithms with a DAG coarsening approach to obtain significantly smaller DAGs, whilst preferably retaining most of the original structure. Our schedulers can then be applied to the coarse graph. The resulting schedule is subsequently pulled back to the original graph to obtain the final schedule. Such a coarsening technique is widely applied in graph partitioning and scheduling tools [KAKS97, PSSS21] and has various advantages for our scheduler: it greatly reduces the size of the graph, improves data locality, and can also help reduce the number of synchronization steps.
In order to produce a valid scheduling problem, the coarsening needs to ensure that it preserves the acyclicity of the DAG. Methods that fulfill this property have been studied in several works before, see, for example, [CLB94, FER+13, HKU+17, ZCL+22] and references therein. In Section 4, we propose the concept of cascades to generalize the coarsening techniques utilized in [CLB94, §4] and [ZCL+22, §IV.B]. We then formally prove that coarsening techniques based on cascades always preserve acyclicity.
1.1.3. Reordering
Besides the scheduling algorithms above, we also apply a matrix reordering step to drastically improve data locality during the SpTRSV computation. Specifically, once the schedule is developed, we symmetrically permute the matrix according to the schedule, ensuring that values computed after each other on the same core are close to each other in this permuted representation. This idea has already been explored by Rothberg–Gupta in the 1990s [RG92], but it has not been applied in modern SpTRSV baselines, which instead try to make use of existing data locality when deriving a schedule.
1.2. Additional related work
An optimization method for parallel SpTRSV execution, orthogonal to what we have mentioned so far, is to break the lower triangular matrix into blocks and apply different algorithms and allocation of cores to each block individually depending on block type and sparsity pattern [AS89, May09, AYU21, YSAU20]. The blocks may be on the diagonal, which corresponds to a smaller instance of (sparse) triangular solve, or completely off the diagonal, which corresponds to a (sparse) matrix-vector multiplication. This has been particularly impactful for GPU implementations [LLH+16, LNL20]. Our scheduling algorithms would naturally fit in as a building block for such optimizations.
Finally, besides the forward-/backward-substitution algorithm, there are also other methods for solving sparse triangular systems, for example inversion. For this method, we mention the memory-optimal algorithms developed in previous works [AS93, PA92].
1.3. Acknowledgements
We would like to thank Weifeng Liu and Yves Baumann for stimulating conversations on this and surrounding topics.
2. Background
2.1. Graph notation
We model our computations as a directed acyclic graph (DAG) $G=(V,E)$, which consists of a set of vertices $V$ and a set of directed edges $E\subseteq V\times V$. For any vertex $v\in V$, the sets of vertices $\{u\,|\,(u,v)\in E\}$ and $\{u\,|\,(v,u)\in E\}$ are called the parents of $v$ and the children of $v$, respectively. The in- and out-degree of $v$ are the number of parents and children of $v$, respectively. The degree of $v$, denoted by ${\rm deg}(v)$, is the sum of its in- and out-degree. If a vertex of the DAG has no parents/children, then it is called a source/sink vertex respectively. The DAG in our model is also complemented by vertex weights $\omega:V\rightarrow\mathbb{Z}_{>0}$ to indicate the compute cost of each operation.
2.2. Problem definition and notation
When solving sparse triangular systems, we are given a triangular matrix $A=(A_{i,j})_{i,j=1,\dots,n}\in\mathbb{R}^{n\times n}$, a dense vector $b=(b_{1},\dots,b_{n})^{T}\in\mathbb{R}^{n}$, and the goal is to solve the equation $Ax=b$ for the vector $x=(x_{1},\dots,x_{n})^{T}\in\mathbb{R}^{n}$. We assume that $A$ is non-singular, such that all its diagonal elements are non-zero. In case of a lower triangular matrix $A$, there is a natural forward-substitution algorithm for the problem, which iterates through the rows of $A$ in order and computes the values of $x$ as $x_{1}=\tfrac{b_{1}}{A_{1,1}}$, $x_{2}=\tfrac{b_{2}-A_{2,1}x_{1}}{A_{2,2}}$, and in general, as
$x_{i}=\frac{1}{A_{i,i}}\left(b_{i}-\sum_{j=1}^{i-1}A_{i,j}x_{j}\right).$ (2.1)
The case of an upper triangular matrix $A$, a backward-substitution algorithm follows symmetrically in the reverse direction.
In the forward-substitution algorithm (2.1), we say that the computation of $x_{i}$ depends on the value of $x_{j}$, for $j<i$, if and only if there is an increasing sequence $j=\ell_{0}<\ell_{1}<\dots<\ell_{m}=i$ such that each entry $A_{\ell_{k-1},\ell_{k}}$ is non-zero, for $k=1,\dots,m$. If there is no dependency between $x_{i}$ and $x_{j}$, the two corresponding operations can be executed in any order, in particular also in parallel. As such, the operations in the algorithm can naturally be represented as a DAG $G=(V,E)$, where $V=\{1,...,n\}$, the vertex $i$ represents the $i$-th row of $A$, and, for any $i,j\in V$, we have a directed edge $(j,i)\in E$ if and only if $A_{i,j}\neq 0$. See Figure 1.1 for an example. To indicate the compute cost of each operation, the weight $\omega(v)$ of each vertex $v\in V$ in the DAG is simply defined as the number of non-zero entries in the corresponding row of the matrix.
The parallel execution of this DAG then directly corresponds to a parallel execution of the SpTSRSV and many previous works found it more convenient to discuss their scheduling methods for this problem using this DAG representation.
The parallel-scheduling problem above can be most fittingly captured in a bulk-synchronous parallel (BSP) model [Val90a] that assumes global synchronization barriers to split the execution into so-called supersteps. This model is also known as the XPRAM model [Val90b]. A schedule in this model assigns each vertex, i.e., the computation of each $x_{i}$, to one of the $k$ available cores and to a given superstep. A valid schedule must fulfill the precedence constraints of the DAG and ensure that we always have a synchronization barrier between computing a value on one core and using it as input on another core.
Definition 2.1.
A parallel schedule of $G$ consist of assignments $\pi:V\rightarrow\{1,...,k\}$ to cores and $\sigma:V\rightarrow\mathbb{Z}_{>0}$ to supersteps, which fulfill the following properties for each $(u,v)\in E$:
  • $\sigma(u)\leqslant\sigma(v)$;
  • if $\pi(u)\neq\pi(v)$, then $\sigma(u)<\sigma(v)$.
The total cost of a schedule is determined by the workload balance within each superstep and the number of synchronization barriers. The original BSP model includes also communication volume in its cost function. For the SpTRSV application, however, the communication happens in parallel to the computation and resolving the synchronizations. Hence, the latter two dominate and dictate the overall execution time. Synchronous methods from previous works apply the same scheduling model, although often without explicitly referring to BSP, XPRAM, or supersteps.
3. Barrier list schedulers
The idea of adapting list schedulers to a setting with communication barriers has been explored recently by Papp et al. [PAKY24] in the context of abstract BSP scheduling. The barrier-list-scheduler prototype in their paper behaves similarly to a classical list scheduler, assigning concrete time steps to the starting and finishing times of each task. Whenever a core $p$ becomes free, it prioritizes assigning a next vertex to this core from those that are only computable on $p$ in the current superstep, since some of its parents were computed on $p$ in the current superstep. When half of the cores become idle (i.e., cannot be assigned new vertices without communication), a synchronization barrier is inserted.
The skeleton of our schedulers is similar to the base idea of this algorithm, but with many of the ingredients improved. In particular, the essence of the algorithm lies in the priority function that selects the next vertex to assign to a free core when there are multiple vertices ready. We present two schedulers with priority functions tailored towards the problem of developing work-balanced schedules with few synchronization steps. Both of our schedulers define this priority function via assigning priority scores to every vertex that is computable in the current superstep. This score may be core-specific as a vertex may be computable on more than one core. Whenever a core becomes free, our algorithms select the vertex with the highest score for the corresponding core as the next vertex to compute on this core.
Data: A vertex-weighted DAG $G=(V,E,\omega)$, a set of cores $P$, and a parameter $\alpha\in\>]0,1]$.
Result: A schedule consisting of processor assignment $\pi:V\to P$ and superstep assignment $\sigma:V\to\mathbb{Z}_{>0}$.
1 ${\rm superstep}\leftarrow 1$
2 ${\rm endStep}\leftarrow\textnormal{{false}}$
3 ${\rm finishTimes}\leftarrow\{0\}$
4 ${\rm free}[p]\leftarrow\textnormal{{true}}$, for all $p\in P$
5 ${\rm prio}[v],{\rm prio}[p,v]\leftarrow$ initialize priority, for all $v\in V$ and $p\in P$
6 ${\rm ready}_{\rm pool},\,{\rm ready}_{\rm all}\leftarrow\text{{sources of }}G$
7 ${\rm ready}_{p}\leftarrow\emptyset$, for all $p\in P$
8 while $\exists$ unassigned vertex do
9       if ${\rm endStep}$ and ${\rm finishTimes}=\emptyset$ then
10             ${\rm ready}_{p}\leftarrow\emptyset$, for all $p\in P$
11             ${\rm ready}_{\rm all}\leftarrow{\rm ready}_{\rm pool}$
12             ${\rm superstep}\leftarrow{\rm superstep}+1$
13             ${\rm endStep}\leftarrow\textnormal{{false}}$
14             ${\rm finishTimes}\leftarrow\{0\}$
15
16      $t\leftarrow$ earliest time from ${\rm finishTimes}$
17       for all vertices $v$ that finish at $t$ do
18             ${\rm free}[\pi(v)]\leftarrow\textnormal{{true}}$
19             for all children $u$ of $v$ do
20                   if all parents of $u$ are finished then
21                         ${\rm ready}_{\rm pool}.{\rm insert}(u)$
22                         if $\forall\,(w,u)\in E:(\pi(w)=\pi(v)\text{ or }\sigma(w)<{\rm superstep})$ then
23                               ${\rm ready}_{\pi(v)}.{\rm insert}(u)$
24
25
26
27
28      if $\neg{\rm endStep}$ or fill-up possible then
29             while $\exists p$ with ${\rm free}[p]=\textnormal{{true}}$ and ${\rm ready}_{p}\cup{\rm ready}_{\rm all}\neq\emptyset$ do
30                   if $\neg{\rm endStep}$ then
31                         $p,\,v\leftarrow$ choose using priority
32
33                   else
34                         $p,\,v\leftarrow$ choose using fill-up strategy
35
36                  delete $v$ from ${\rm ready}_{\rm pool}$, ${\rm ready}_{p}$, and ${\rm ready}_{\rm all}$
37                   $\pi(v)\leftarrow p$
38                   $\sigma(v)\leftarrow{\rm superstep}$
39                   ${\rm finishTimes}.{\rm insert}(t+w(v))$
40                   ${\rm free}[p]\leftarrow\textnormal{{false}}$
                   update priority in $v$’s neighborhood
                    // only in Locking
41
42
43      if $\alpha$ fraction of cores idle and ${\rm ready}_{\rm pool}$ large enough  then
44             ${\rm endStep}\leftarrow\textnormal{{true}}$
45
46      ${\rm finishTimes}.{\rm remove}(t)$
47
48return $(\pi,\sigma)$
Algorithm 3.1 Skeleton of our scheduling algorithms
  • In the $p$-ivotal path scheduler, we attempt to prioritize vertices that have long outgoing paths and/or a high number of successors in the (vertex-weighted) DAG $G=(V,E,\omega)$. For this, we use a static scoring function. The score of each vertex $v\in V$ is computed recursively using the parameter $p\geq 1$:
    ${\rm prio}[v]=\omega(v)+\left(\sum_{u\in{\rm Children}(v)}{\rm prio}[u]^{p}\right)^{\frac{1}{p}}.$ (3.1)
    Note that choosing $p\to\infty$ is just the longest path to a sink vertex excluding communication, which is sometimes referred to as bottom (vertex) distance. In our implementation, we have simply chosen $p=2$. We remark that one needs to be wary of numerical instability (overflow) in the implementation of this priority.
  • In the Locking scheduler, we aim to specifically avoid vertices that would lock out some other vertex from the current superstep due to a communication requirement. In particular, the assignment of vertex $v$ to a core $p$ gets a penalty value of $\ell$ if there are $\ell$ yet-uncomputed children of $v$ that already have parents computed on (at most one) core other than $p$ in the current superstep, i.e., assigning $v$ to $p$ would imply that these $\ell$ vertices can be computed in the next superstep at the earliest. This strategy aims to increase the number of vertices we can compute before having to insert a synchronization barrier. Besides this, to the score of each vertex, we also add the length of the longest outgoing path normalized to a small range $[0,20]$ as a default initialization value, in order to prioritize progress on the critical paths in general. Note that in this scheduler, whenever a vertex is assigned to a core, the priority values on nearby vertices need to be updated to correctly reflect the new number of vertices that would be locked out by each potential assignment. It is thus of dynamic nature.
Our algorithms use dynamically updated priority queues to ensure that the scoring only imposes a minimal overhead on the scheduling.
Besides this, our schedulers are also extended with a fill-up strategy: whenever the algorithm decides that a synchronization barrier is required at time $t$, we specifically check for each core $p$ whether some lower-weight vertices could still fit into the current computation phase on $p$ without delaying the synchronization barrier, and compute as many of these vertices as we can in order of their priority. This is especially useful for DAGs obtained after coarsening, where the compute weights of vertices can differ significantly.
Finally, our schedulers also use a much more sophisticated rule to decide when to close the current superstep. The fraction $\alpha$ of cores that need to be idle to insert a barrier is changed to a parameter $\alpha$, chosen between $0.2$ and $0.4$ for our algorithms. Moreover, our schedulers ensure that even when this threshold is met, our algorithm only inserts a barrier when the current part of the DAG indeed allows for enough parallelization in order to provide work for idle cores.
Altogether, the run time of both of our schedulers is just above linear in the size of the input graph, i.e., in the number of non-zero entries in the matrix. In particular, the worst-case theoretical time complexity is $O(|V|\log|V|+|E|)$ for $p$-ivotal path, and $O(k|V|\log|V|\cdot\sum_{v\in V}{\rm deg}(v)^{2}+|E|)$ for Locking, where $k$ is the number of cores. In practice, our experiments show that the run time of the two schedulers is similar in most cases, and also comparable to that of the HDagg baseline, cf. Section 7.7. In terms of space complexity, $p$-ivotal path requires $O(|V|+|E|)$, whilst Locking requires $O(k|V|+|E|)$ memory.
We provide a detailed pseudocode for our schedulers in Algorithm 3.1, and note:
  • The $\text{ready}_{\text{all}}$ and $\text{ready}_{p}$ structures defined in Lines 3.1, 3.1 of Algorithm 3.1 are priority queues that are always sorted according to the priority values ${\rm prio}[v]$ and ${\rm prio}[p,v]$ of the contained vertices $v$, respectively. It is due to the insertions/deletions from these priority queues that the run time of the schedulers is not strictly linear.
  • In Line 3.1 of Algorithm 3.1, the last condition is specifically that
    $|{\rm ready}_{\rm pool}|\geqslant\min\left(1.2\cdot{\rm busy},\,{\rm busy}+\tfrac{1}{2}\cdot{\rm idle}\right),$ (3.2)
    with ‘${\rm idle}$’ and ‘${\rm busy}$’ denoting the current number of idle and busy cores, respectively. The condition guarantees that inserting a synchronization barrier indeed allows us to employ more cores. Intuitively, the first and second term in the minimum ensure that the increase in parallelism is significant in the cases when ‘${\rm busy}$’ is a small and large value, respectively.
  • Once we decide to insert a barrier (the endStep variable is set), the finishing time $t_{\rm end}$ of the current superstep is finalized as the latest finishing time of vertices that are currently being computed, i.e., the latest entry in finishTimes. Then, until the superstep ends, the fill-up strategy is used: we only start computing a vertex $v$ on a core $p$ if this extra computation is still finished by $t_{\rm end}$ at the latest. In Line 3.1, for each core $p$, we again select $v$ as the highest-priority vertex in $\text{ready}_{p}$ or $\text{ready}_{\text{all}}$ that satisfies this condition.
  • In Line 3.1, the $p$-ivotal path scheduler initializes all vertices of the DAG in linear time using Equation (3.1), via a dynamic programming approach that follows a reverse topological ordering.
  • In contrast to this, in Line 3.1, the Locking scheduler also uses dynamic programming to compute the longest outgoing path from each vertex, then normalizes this to the range $[0,20]$, and uses it as a default priority value. Then, whenever a new vertex is added to the priority queues (in Lines 3.1, 3.1, or 3.1), the rest of the priority function (the number of vertices locked out from the superstep by the assignment) is computed and added to this base priority value.
  • In the Locking scheduler, Line 3.1 considers all children $u$ of $v$ that are not yet locked out of the current superstep. If no parent of $u$ is in the current superstep yet, then for all cores except $p$, the (ready) parents of $u$ get an extra unit of penalty for locking $u$ out from the superstep. If $u$ already has parents on (only one) other core than $p$ in the current superstep, then it is our current assignment that locks $u$ out of the superstep, so all the (ready) parents of $u$ lose a unit of penalty.
4. Acyclicity-preserving graph coarsening
Previous works discuss several ways to partition a DAG into clusters such that this coarsened graph remains acyclic, although often without a formal proof of this property. In a further generalization of earlier methods from Cong et al. [CLB94, §4] and Zarebavani et al. [ZCL+22, §IV.B], we now introduce the concept of cascades and we prove that coarsening a DAG along such cascades is still guaranteed to preserve acyclicity. This is presented in Section 4.1. In Section 4.2, we describe the graph coarsening algorithm used in our scheduling algorithms.
4.1. Cascades
We begin with some formal definitions. Thereafter, we prove Proposition 4.3, demonstrating the utility of cascades for coarsening DAGs.
Definition 4.1.
Let $G=(V,E)$ be a directed graph and $P$ a partition of $V$. We define the coarsened graph of $G$ along $P$ as the graph $(V^{\prime},E^{\prime})$, where $V^{\prime}=P$, i.e., the vertices are the parts of the partition $P$, and for $U^{\prime},W^{\prime}\in V^{\prime}$ we have that $(U^{\prime},W^{\prime})\in E^{\prime}$ if and only if $U^{\prime}\neq W^{\prime}$ and $\exists(u,w)\in E$ such that $u\in U^{\prime}$ and $w\in W^{\prime}$. We denote the coarsened graph of $G$ along $P$ by $G/\mkern-5.5mu/P$.
In other words, the coarsened graph $G/\mkern-5.5mu/P$ is the graph $G$ quotiented by the equivalence relation induced by $P$ with self-loops removed. The definition is easily extended to vertex-weighted graphs, where the weight of a part $U\in P$ is given as the sum the weights of its elements: $\omega(U)=\sum_{u\in U}\omega(u)$.
Definition 4.2.
Let $G=(V,E)$ be a directed graph. We call a subset of vertices $U\subseteq V$ a cascade if and only if for every vertex $v\in U$ with an incoming cut edge, that is $(w,v)\in E$ such that $w\not\in U$, and for every vertex $u\in U$ with an outgoing cut edge, that is $(u,w)\in E$ such that $w\not\in U$, there is a (possibly trivial) directed walk from $v$ to $u$ in $G$.
Proposition 4.3.
Let $G=(V,E)$ be a directed acyclic graph and $P$ a partition of $V$ such that each set $U\in P$ is a cascade. Then, the coarsened graph $G/\mkern-5.5mu/P$ of $G$ along $P$ is acyclic.
Proof.
We will show that any directed walk in $G/\mkern-5.5mu/P$ can be elevated to a directed walk in $G$. Therefore, the existence of directed cycles in $G/\mkern-5.5mu/P$ implies the existence of directed cycles in $G$.
We lift a walk from $G/\mkern-5.5mu/P$ by mapping each edge to an arbitrary representative in $E$, whose endpoints necessarily lie in disjoint sets of the partition $P$ as $G/\mkern-5.5mu/P$ does not contain any self-loops, and connecting the endpoints via the directed walks guaranteed by the defining property of cascades. ∎
4.2. Algorithm
In our graph coarsening algorithm, we do not make use of the full strength of Proposition 4.3. Instead, we use a subcategory of cascades, which can be found efficiently. We call them funnels, though they have been previously described under the name fanout-free cone [CLB94, §4]. Since the latter reference does not include an algorithm with a complexity analysis, we include them here in Algorithm 4.1.
Definition 4.4.
Let $G=(V,E)$ be a directed acyclic graph. We call a subset of vertices $U\subseteq V$ an in-funnel if and only if $U$ is a cascade and there is at most one vertex $u\in U$ with an outgoing cut edge, that is $(u,w)\in E$ such that $w\not\in U$.
We analogously define an out-funnel.
// Main algorithm
Data: A directed acyclic graph $G=(V,E)$.
Result: A partition $P$ such that every set $U\in P$ is an in-funnel.
1 Algorithm ${\rm Funnel}(G)$:
2       ${\rm Partition}\leftarrow\emptyset$
3       ${\rm Visited}[v]\leftarrow\textnormal{{false}},\,\forall v\in V$
4       for $v\in V$ in reverse topological order do
5             if ${\rm Visited}[v]$ then continue
6             $U\leftarrow\{v\}$
             $(U,\_)\leftarrow{\rm FunnelDFS}(G,U,v,\emptyset)$
              // see subroutine, Line 4.1
7             for $u\in U$ do
8                   ${\rm Visited}[u]\leftarrow$ true
9
10            ${\rm Partition}.{\rm insert}(U)$
11
12      return Partition
13
// Depth-first search invoked by the main algorithm
Data: A directed acyclic graph $G=(V,E)$, a non-empty subset $U\subseteq V$ of vertices, a vertex $u\in U$, and a hash map ${\rm IncludedChildren}$.
Result: A non-empty subset $W\subseteq V$ containing $U$ and a hash map ${\rm IncludedChildren}$.
14 Subroutine ${\rm FunnelDFS}(G,U,u,{\rm IncludedChildren})$ :
15       for $v\in{\rm Parents}(u)$  do
16             if ${\rm IncludedChildren.find}(v)$ then
17                  ${\rm IncludedChildren}[v]\leftarrow{\rm IncludedChildren}[v]+1$
18             else
19                  ${\rm IncludedChildren}[v]\leftarrow 1$
20            if ${\rm IncludedChildren}[v]={\rm OutDegree}(v)$  then
21                   $U.{\rm insert}(v)$
22                   $(U,{\rm IncludedChildren})\leftarrow{\rm FunnelDFS}(G,U,v,{\rm IncludedChildren})$
23
24
25      return $(U,{\rm IncludedChildren})$
26
27
Algorithm 4.1 In-funnel graph coarsening.
The time complexity of the topological sort is $O(|V|+|E|)$ [Kah62] and its space complexity is $O(|V|)$. In order to bound the time complexity for the remaining part, we note that each parent vertex $v$ in Line 4.1 is visited at most as many times as its out-degree, leading to an overall complexity of $O(|V|+|E|)$. The space complexity is easily seen to be $O(|V|)$.
In practice, before applying this graph coarsening, we remove some transitive edges from $G$ as this increases the likelihood of finding larger components. A complete transitive reduction is slow, though there are faster approximate transitive reductions, such as the ‘remove all long edges in triangles’-algorithm [PSSD14, §2.3] with a time complexity of $O(\sum_{v\in V}{\rm deg}(v)^{2})$. This algorithm may be terminated early if a faster runtime is desired. In our implementation, we run the full algorithm.
In our implementation, we also add a size/weight constraint on each component of the partition to Algorithm 4.1 as otherwise a graph with only one sink vertex would be coarsened into a graph with only one vertex.
Finally, we remark that we have observed experimentally that coarsening along in-funnels works better than out-funnels. Whether this has something to do with the graphs in the data set or there is an underlying conceptual reason for this is left as an open question.
5. Reordering for locality
Our algorithms already account for two of the most important factors in synchronous scheduling: work balance and the number of synchronization barriers. However, another major aspect that greatly influences the efficiency of a parallel SpTRSV execution is data locality, i.e., the number of required values that are already available in cache.
In order to address this, we apply a separate reordering step to ensure that vertices which are computed together are also preferably stored together. The main idea of this approach has already been considered before, cf. [RG92], but has not found its way into modern baselines. In particular, we consider a reordering (relabeling) the vertices of the input DAG based on the partitioning we developed, where we iterate through the supersteps in order, and within each superstep, we iterate through the cores in order. That is, we first start with the vertices $v$ with $\pi(v)=1$, $\sigma(v)=1$, then the vertices $v$ with $\pi(v)=2$, $\sigma(v)=1$, and so on, up to the vertices $v$ with $\pi(v)=k$, $\sigma(v)=1$, followed by vertices $v$ with $\pi(v)=1$, $\sigma(v)=2$, etc. Within a given core-superstep combination, we go through the vertices in the original order (which gives a topological ordering of the induced subDAG). We then symmetrically permute the input matrix and permute the right-hand-side vector of the SpTRSV problem accordingly. Note that since the permutation provides a valid topological ordering of the vertices of the DAG, the resulting matrix is still lower triangular, resulting in an equivalent (symmetrically permuted) formulation of the SpTRSV problem.
We then execute the SpTRSV computation on the permuted problem, following our schedule, which ensures that vertices computed on the same core in the same superstep are stored close to each other, thus greatly improving locality during the computation.
6. Experimental setup
In this section, we present the experimental setup for the evaluation of our scheduling algorithms. Our implementations are available in the OneStopParallel repository [BLPS24] on Github.
6.1. Methodology
For the evaluation, we used a standard SpTRSV implementation which iterates through the rows of the matrix which was stored in compressed sparse row (CSR) format [TW67]. The algorithm was parallelized using the OpenMP library with the flags OMP_PROC_BIND=close and OMP_PLACES=cores.
We measured one hundred times the time it takes for a single SpTRSV execution using the chrono high-resolution clock. The measurements were taken whilst the system was ‘hot’, meaning an untimed execution precedes the timed executions. Between each SpTRSV execution, the right-hand-side vector $b$ was reset to all ones. The experiments were repeated for each scheduling algorithm, matrix in the data set, and CPU architecture type. The latter two are described in more detail in Section 6.2 and Section 6.3, respectively. If the interquartile range of the measurements corresponding to a scheduling method was too large, we rejected and re-ran all experiments on the same matrix and processor configuration.
The experiments for the schedulers HDagg and SpMP were carried out in the sympiler framework [CKSD17, Che22] as in [ZCL+22] with only minor adjustments to adhere to the aforementioned setup. All remaining schedulers were tested in our own framework.
All scheduling algorithms are implemented in C++ and were compiled with GCC using the optimization flag -O3.
6.2. Data sets
For the experiments, we used matrices from several data sets. The main data set is a sample from the SuiteSparse Matrix Collection [DH11], which constitutes a diverse set of matrices from a wide range of applications and was used in previous studies [ZCL+22]. This data set is complemented with two randomly generated ones: uniformly random, i.e., Erdős–Rényi matrices [ER59], and random with a bias towards the diagonal. The former are easier to parallelize as they have few (and thus large) wavefronts [HKSL14] and the latter are designed to be harder to parallelize.
A useful general metric to understand the parallelizability of an SpTRSV execution is the average wavefront size, which can be calculated from the DAG representation by dividing the number of vertices by the length of the longest path. This metric is also indicated for each matrix in the overview of the data sets in the appendices.
6.2.1. SuiteSparse
From the SuiteSparse Matrix Collection [DH11], we used the lower triangular part of all the sparse real symmetric positive definite matrices. Out of those, we further restricted ourselves to large matrices with enough available parallelism, meaning
  • the number of floating point operations [^1] The number of floating point operations is equal to twice the number of non-zeros minus the dimension of the matrix. is at least $2$ million, and
  • the average wavefront size is at least $44$, twice the number of cores utilized in the experiments.
We furthermore removed matrices from the data set which had the same sparsity pattern. An overview over some statistics of the matrices may be found in Table LABEL:table:florida-graph-table.
Remark 6.1.
Zarebavani et al. [ZCL+22] also use the real symmetric positive definite matrices from the SuiteSparse Matrix Collection for their data set. Their data set, however, differs from ours insofar as they first symmetrically permute the sparse symmetric matrix using a fill-reducing method of METIS [KK98] and only then take the lower triangular part. In general, this results in non-equivalent SpTRSV problems. As a by-product of this extra step, the matrices in their data set have significantly higher (14×14$\times$14 × geometric-mean) average wavefront size than the lower triangular part of the original matrices, compare Table LABEL:table:hdagg-metis-graph-table with Table LABEL:table:florida-graph-table. As a result, the performance characteristics of HDagg and SpMP reported in their paper differ significantly from our experiments.
We mention that in some contexts, such as with the incomplete Cholesky preconditioned conjugate gradient method for sparse symmetric solve, one is allowed to first symmetrically permute to increase the average wavefront size.
6.2.2. Erdős–Rényi
[…]
We generated thirty $N\times N$ matrices of this type with $N=$ 100,000 and $p=10^{-4},5\cdot 10^{-4},2\cdot 10^{-3}$, ten of each given probability. An overview over some statistics of the matrices may be found in Table LABEL:table:erdos-renyi-graph-table.
6.2.3. Narrow bandwidth
Unlike the Erdős–Rényi random matrices, we let the lower triangular matrix entry $(i,j)$, with $i>j$, being independently non-zero with probability $p\cdot\exp((1+j-i)/B)$, moving the non-zero entries closer to the diagonal. The entry values were chosen as in Section 6.2.2.
We generated thirty $N\times N$ matrices of this type with $N=$ 100,000 and $(p,B)=(0.14,10)$, $(0.05,20)$, $(0.03,42)$, ten for each pair $(p,B)$. An overview over some statistics of the matrices may be found in Table LABEL:table:random-bandwidth-graph-table.
6.3. CPU architectures
The CPU architectures used for the experiments were x86 and ARM. The precise model and some specifications are given, respectively, as follows:
  • Intel Xeon Gold 6238T processor (x86), with 192 GB memory and theoretical peak memory throughput of 140.8 GB/s and 22 cores on a single socket;
  • AMD EPYC 7763 processor (x86), with 1024 GB memory and theoretical peak memory throughput of 204.8 GB/s and 64 cores on a single socket;
  • Huawei Kunpeng 920-4826 (Hi1620) processor (ARM), with 512 GB memory and theoretical peak memory throughput of 187.7 GB/s and 48 cores on a single socket.
7. Evaluation
7.1. Overall performance
We present speed-ups of the forward-/backward-substitution algorithm based on parallel schedules compared to serial execution. The schedules of our proposed algorithms are benchmarked against those produced by the baseline methods, SpMP [PSSD14] and HDagg [ZCL+22]. The results, aggregated over the instances from the respective data set using the geometric mean of all pairs of runs, are displayed in Table 7.1. All experiments were conducted on the Intel x86 machine utilizing 22 cores.
On our main data set, SuiteSparse, the schedules generated by our Funnel Locking algorithm achieved a geometric-mean speed-up of 1.45×1.45$\times$1.45 × compared to SpMP and 3.24×3.24$\times$3.24 × compared to HDagg. The data shows a similarly large improvement on the Erdős–Rényi data set. Here, the $p$-ivotal path algorithm achieved the best results, with a speed-up of 1.62×1.62$\times$1.62 × over SpMP and 1.87×1.87$\times$1.87 × over HDagg.
The narrow bandwidth graphs paint a slightly different picture. Here, the best schedule was produced by SpMP, with our algorithms falling behind. This is because these matrices allow for significantly less parallelizaton by design; such a setting is not a good fit for our schedulers, which aim to utilize most of the cores. In particular, Table 7.1 shows that when only fewer cores are available, our algorithms again outperform the baselines on this data set.
{tblr}
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set & (Funnel) Locking (Funnel) $p$-ivotal path SpMP HDagg
SuiteSparse 10.709.35 10.449.27 7.39 3.30
Erdős–Rényi 14.0814.09 14.9215.00 9.23 7.96
Narrow bandw. 2.792.40 3.252.53 4.37 1.09
Table 7.1. Geometric mean of speed-ups over serial execution of $p$-ivotal path and Locking with/without Funnel coarsening, compared to the baselines SpMP and HDagg on the Intel x86 machine using 22 cores taken over the data sets described in Section 6.2.
{tblr}
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set (Funnel) Locking (Funnel) $p$-ivotal path SpMP HDagg
Narrow bandw. 4.16 / 3.87 4.60 / 4.36 3.67 1.96
Table 7.2. Geometric mean of speed-ups over serial execution of $p$-ivotal path and Locking with/without Funnel coarsening, compared to the baselines SpMP and HDagg on the Intel x86 machine using 8 cores taken over the narrow bandwidth data set.
We also include a performance profile [DM02] based on the data generated from the SuiteSparse data set in Figure 7.1. The closer the line is to the top left corner, the better and more consistent the algorithm is across the data set. This shows that our algorithms are not only faster in execution time on average but are so throughout the diverse SuiteSparse data set.
Figure 7.1. Performance profiles of various algorithms on the SuiteSparse data set evaluated on the Intel x86 machine using 22 cores. The x-axis represents the threshold and the y-axis is the proportion of runs that are within the threshold times fastest SpTRSV run on the respective matrix.
7.2. Fewer synchronization barriers
The result in Table 7.1 show that our schedulers can significantly outperform the synchronous state-of-the-art HDagg. A further analysis shows that this is in part due to a substantial reduction in the number of synchronization barriers required during execution, whilst still maintaining a good work balance. In particular, Table 7.2 shows the number of synchronization barriers relative to the number of wavefronts in our algorithms and HDagg. The data indicates a large, up to 14×14$\times$14 ×, reduction of number of synchronization barriers compared to the number of wavefronts. This is a reduction of up to $\times1$ compared to HDagg, which explains the significant speed-ups achieved by our methods.
{tblr}
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, guard, rowhead = 1, rowfoot = 0, Data set (Funnel) Locking (Funnel) $p$-ivotal path HDagg
SuiteSparse 14.00 / 9.63 13.26 / 9.76 1.24
Erdős–Rényi 2.75 / 2.74 3.29 / 3.26 1.25
Narrow bandw. 4.53 / 4.01 6.08 / 4.01 1.10
Table 7.3. Geometric mean of the reduction of the number of synchronization barriers relative to the number of wavefronts of the matrix.
7.3. Comparison of our algorithms
To compare our techniques and algorithms, we compute the geometric-mean speed-ups over Serial. We present average and interquartile ranges in Figure 7.2. The data shows that the Locking algorithms slightly outperform the $p$-ivotal path ones on the SuiteSparse data set, while the $p$-ivotal path algorithms have an edge on the Erdős–Rényi and the narrow bandwidth data sets. We observe that applying our coarsening algorithm, Funnel, greatly improves results on the SuiteSparse and the narrow bandwidth data sets. However, on the less structured matrices in the Erdős–Rényi data set, coarsening shows no performance gain.
(a) SuiteSparse
(b) Erdős–Rényi
(c) Narrow bandwidth
Figure 7.2. Geometric mean and interquartile ranges of speed-ups of (Funnel) $p$-ivotal path and (Funnel) Locking relative to Serial grouped according to data set. Experiments were conducted on the Intel x86 machine using 22 cores.
7.4. Impact of reordering
We separately analyze the impact of the reordering step on the performance. Table 7.4 compares the speed-ups achieved by our algorithms with and without the reordering component from Section 5. The numbers show that reordering is indeed a valuable ingredient of our schedulers. The data also confirms that even without the reordering, the algorithms still notably outperform HDagg, which is the current state-of-the-art synchronous baseline, cf. Table 7.1. Note that the narrow bandwidth random graphs are close to sequential by design, allowing limited parallelization, and thus any valid schedule here already exhibits good locality without reordering; as such, the reordering technique is not beneficial in this case. How to best reorder to improve locality for SpTRSV is research topic that deserves further attention.
{tblr}
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, rowhead = 1, rowfoot = 0, Data set Funnel Locking Locking Funnel $p$-ivotal p. $p$-ivotal p.
SuiteSparse 10.70 / 7.86 9.35 / 7.36 10.44 / 7.42 9.27 / 7.03
Erdős–Rényi 14.08 / 8.91 14.09 / 8.95 14.92 / 9.10 15.00 / 9.10
Narrow bandw. 2.79 / 3.19 2.40 / 2.64 3.25 / 3.20 2.53 / 2.69
Table 7.4. Geometric mean of speed-ups relative to Serial of our algorithms with/without permuting the matrix data according to the computed schedule. Experiments were conducted on the Intel x86 machine using 22 cores.
7.5. Performance across different architectures
We show the performance gains of our algorithms over the different processors and architectures in Table 7.5. The data confirms that our algorithms consistently outperform the baselines across all considered architectures. We note that the improvement relative to Serial can be in a significantly different range due to the properties of the distinct architectures. SpMP is omitted for the ARM architecture because its implementation is x86-specific.
{tblr}
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, rowhead = 1, rowfoot = 0, Machine (Funnel) Locking (Funnel) $p$-ivotal path SpMP HDagg
Intel x86 10.70 / 9.35 10.44 / 9.27 7.39 3.30
AMD x86 5.99 / 5.40 5.80 / 5.42 4.39 2.12
Huawei ARM 9.30 / 8.89 9.07 / 8.80 n/a 2.03
Table 7.5. Geometric mean speed-ups relative to Serial of our algorithms over different machines and processor architectures. Experiments were conducted using 22 cores on the SuiteSparse data set.
7.6. Scaling with the number of cores
Another natural question is how our algorithms scale with a growing number of cores. To examine this, we illustrate the speed-ups (over serial execution) for different numbers of cores in Table 7.6. We note that this experiment was conducted on the AMD x86 machine as it has $64$ available cores on a single socket.
{tblr}
colspec = Q[l,m] || Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] | Q[c,m] , row1 = font=, rowhead = 1, rowfoot = 0, Algorithm 2 4 8 16 24 32 40 48 56 64
Funnel Locking 1.86 2.92 3.40 5.24 6.19 6.72 6.92 7.02 6.87 6.65
Funnel $p$-ivotal path 1.85 2.89 3.31 5.11 6.04 6.45 6.62 6.76 6.63 6.52
Table 7.6. Geometric mean of speed-ups relative to Serial of Funnel $p$-ivotal path and Funnel Locking for different number of cores on the AMD x86 machine taken over the SuiteSparse data set.
As one sees, additional cores have diminished or negative returns at the higher end of number of cores. A reason for this is the average wavefront size which is a proxy for the amount of parallelism available. If we split the SuiteSparse data set into groups according to their average wavefront size, we see that these groups scale to different number of cores, see Figure 7.3. This shows that our algorithms do scale if the matrices allow for it and that the number of cores is an important parameter.
Figure 7.3. Geometric mean speed-ups of Funnel $p$-ivotal path for different number of cores on the AMD x86 machine taken over the SuiteSparse data set categorized by average wavefront size.
7.7. Amortization of scheduling time
In this final section, we consider the gain of the different scheduling algorithms when the scheduling time is taken into account. We measure the amortization threshold as the ratio of the time it took to compute the schedule to the difference in execution times between the serial execution and the parallel execution using the respective schedule [^3] If the parallel execution is slower than the serial one, then the amortization threshold is defined as +$\infty$+ ∞.. The same metric was considered by Zarebavani et. al. [ZCL+22, §V.B] and expresses how often the schedule needs to be reused in order to justify the time spent on computing it. Table 7.7 presents the amortization threshold for several scheduling algorithms with the 25th percentile, median, and 75th percentile values shown for each algorithm.
{tblr}
colspec = Q[l,m] || Q[si=table-format=3.2, table-number-alignment=center,c] | Q[si=table-format=4.2, table-number-alignment=center,c] | Q[si=table-format=4.2, table-number-alignment=center,c] , row1 = font=, guard, rowhead = 1, rowfoot = 0, cell62 = font=, cell63 = font=, cell64 = font=, Algorithm 25th percentile Median 75th percentile
Funnel Locking 740.96 870.26 1158.46
Funnel $p$-ivotal path 731.91 867.67 1047.27
Locking 83.58 111.36 176.16
$p$-ivotal path 52.14 63.92 120.73
SpMP 4.91 6.44 8.93
HDagg 489.88 1306.06 2065.96
Table 7.7. Amortization threshold of several scheduling algorithms on the SuiteSparse data set with the 25th percentile, median, and 75th percentile values shown for each algorithm. The data was collected on the Intel x86 machine using 22 cores.
A noticeable gap exists between the funneled and un-funneled versions of both the $p$-ivotal path and Locking algorithms. The funneled versions involve coarsening, a process that is computationally expensive and results in higher scheduling times. However, this added cost is justified when schedules are reused more than 1000 times in magnitude. In such scenarios, the coarsening pays off by reducing execution time, leading to a greater overall benefit in repeated parallel execution performance.
SpMP consistently exhibits low scheduling times across all percentiles, resulting in exceptional amortization threshold. This reflects the strong engineering behind SpMP, which minimizes scheduling overhead compared to the other algorithms. Our scheduling algorithms are research prototypes and single threaded, and as such, their run time could likely be reduced with further engineering. For example, the matrices may be split into blocks, c.f. Section 1.2, and our algorithms can be applied to each triangular block in parallel. The resulting block schedules are then easily combined into one schedule for the whole matrix. In any case, in scenarios where schedules are reused often enough, our proposed algorithms in their current implementation already demonstrate significant reduction in overall execution time.
8. Conclusion and future directions
The results show that our barrier list schedulers indeed significantly speed up the parallel SpTRSV kernel, reducing the execution time by a 1.45×1.45$\times$1.45 × geometric-mean factor compared to SpMP and 3.24×3.24$\times$3.24 × compared to HDagg on the SuiteSparse benchmark. The data also shows that the three main components of our approach (algorithms, coarsening, and reordering) indeed all contribute to the speed-up, and that the improvement is consistent over various matrix types and architectures.
Future work may consider the adaptation our algorithms to non-uniform memory access (NUMA) architectures. In particular, the AMD x86 data in Section 7.6 confirms that our algorithms scale well to a high number of cores. However, when solving SpTRSV on highly NUMA architectures, we expect our parallel execution schedules to be less effective. It is an interesting question whether one can develop scheduling algorithms that can also efficiently adapt to NUMA, for example, by considering non-uniform bandwidth or latency.
Another promising direction for future work is to combine our barrier list scheduling algorithms with other approaches that proved successful for SpTRSV in the past. For instance, one could seek to merge our algorithms with the block decomposition techniques described in Section 1.2, or to adapt them to a semi-asynchronous setting as in SpMP, in order to allow for a more flexible parallel execution. These methods could allow for further speed-ups on top of our current results.
Appendix A Tables of matrices
We also include some basic statistics of the matrices used in the experiments, cf. Section 6.2.
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics from SuiteSparse Matrix Collection [DH11] used for the evaluation. The average wavefront size has been rounded down., label = table:florida-graph-table, ] colspec = Q[l,m] | Q[r,m] | Q[r,m] | Q[r,m] , row1 = font=, rowhead = 1, rowfoot = 0, Matrix Dimension #Non-zeros Avg. wavefront
af_0_k101 503,625 9,027,150 74
af_shell7 504,855 9,046,865 135
apache2 715,176 2,766,523 1,077
audikw_1 943,695 39,297,771 203
bmw7st_1 141,347 3,740,507 199
bmwcra_1 148,770 5,396,386 204
bone010 986,703 36,326,514 470
boneS01 127,224 3,421,188 156
boneS10 914,898 28,191,660 386
Bump_2911 2,911,419 65,320,659 283
bundle_adj 513,351 10,360,701 57,039
consph 83,334 3,046,907 139
Dubcova3 146,689 1,891,669 44
ecology2 999,999 2,997,995 500
Emilia_923 923,136 20,964,171 176
Fault_639 638,802 14,626,683 143
Flan_1565 1,564,794 59,485,419 200
G3_circuit 1,585,478 4,623,152 611
Geo_1438 1,437,960 32,297,325 246
hood 220,542 5,494,489 365
Hook_1498 1,498,023 31,207,734 95
inline_1 503,712 18,660,027 287
ldoor 952,203 23,737,339 141
msdoor 415,863 10,328,399 59
offshore 259,789 2,251,231 75
parabolic_fem 525,825 2,100,225 75,117
PFlow_742 742,793 18,940,627 118
Queen_4147 4,147,110 166,823,197 342
s3dkt3m2 90,449 1,921,955 60
Serena 1,391,349 32,961,525 298
shipsec1 140,874 3,977,139 67
StocF-1465 1,465,137 11,235,263 487
thermal2 1,228,045 4,904,179 991
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics in the Erdős–Rényi data set used for the evaluation. The average wavefront size has been rounded down., label = table:erdos-renyi-graph-table, ] colspec = l | r | r | r , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeroes Avg. wavefront
ErdosRenyi_100k_19m_A 100,000 19,999,021 109
ErdosRenyi_100k_19m_B 100,000 19,998,182 109
ErdosRenyi_100k_19m_C 100,000 19,997,897 107
ErdosRenyi_100k_19m_D 100,000 19,995,405 106
ErdosRenyi_100k_19m_E 100,000 19,994,516 107
ErdosRenyi_100k_19m_G 100,000 19,989,535 106
ErdosRenyi_100k_19m_H 100,000 19,999,989 110
ErdosRenyi_100k_1m_A 100,000 1,001,528 1,785
ErdosRenyi_100k_1m_B 100,000 1,000,452 1,818
ErdosRenyi_100k_1m_C 100,000 1,000,315 1,818
ErdosRenyi_100k_1m_E 100,000 1,000,044 1,666
ErdosRenyi_100k_1m_F 100,000 1,000,406 1,785
ErdosRenyi_100k_1m_G 100,000 1,001,171 1,724
ErdosRenyi_100k_1m_H 100,000 1,001,551 1,886
ErdosRenyi_100k_1m_I 100,000 1,000,237 1,639
ErdosRenyi_100k_1m_J 100,000 1,001,533 1,851
ErdosRenyi_100k_20m_F 100,000 20,001,732 107
ErdosRenyi_100k_20m_I 100,000 20,006,442 109
ErdosRenyi_100k_20m_J 100,000 20,003,479 109
ErdosRenyi_100k_4m_A 100,000 4,998,205 395
ErdosRenyi_100k_4m_C 100,000 4,999,271 398
ErdosRenyi_100k_4m_G 100,000 4,999,358 401
ErdosRenyi_100k_4m_J 100,000 4,996,501 414
ErdosRenyi_100k_5m_B 100,000 5,006,107 411
ErdosRenyi_100k_5m_D 100,000 5,001,575 404
ErdosRenyi_100k_5m_E 100,000 5,004,251 400
ErdosRenyi_100k_5m_F 100,000 5,002,190 400
ErdosRenyi_100k_5m_H 100,000 5,000,573 409
ErdosRenyi_100k_5m_I 100,000 5,001,846 400
ErdosRenyi_100k_999k_D 100,000 999,915 1,818
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics in the narrow bandwidth data set used for the evaluation. The average wavefront size has been rounded down., label = table:random-bandwidth-graph-table, ] colspec = l | r | r | r , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeroes Avg. wavefront
RandomBand_p14_b10_100k_146k_A 100,000 146,565 87
RandomBand_p14_b10_100k_146k_B 100,000 146,328 115
RandomBand_p14_b10_100k_146k_D 100,000 146,972 73
RandomBand_p14_b10_100k_146k_F 100,000 146,855 111
RandomBand_p14_b10_100k_146k_J 100,000 146,781 105
RandomBand_p14_b10_100k_147k_C 100,000 147,201 61
RandomBand_p14_b10_100k_147k_E 100,000 147,369 73
RandomBand_p14_b10_100k_147k_G 100,000 147,350 132
RandomBand_p14_b10_100k_147k_H 100,000 147,412 85
RandomBand_p14_b10_100k_147k_I 100,000 147,132 132
RandomBand_p3_b42_100k_127k_A 100,000 127,045 46
RandomBand_p3_b42_100k_127k_B 100,000 127,019 55
RandomBand_p3_b42_100k_127k_C 100,000 127,708 29
RandomBand_p3_b42_100k_127k_D 100,000 127,341 45
RandomBand_p3_b42_100k_127k_E 100,000 127,569 67
RandomBand_p3_b42_100k_127k_F 100,000 127,137 47
RandomBand_p3_b42_100k_127k_G 100,000 127,774 52
RandomBand_p3_b42_100k_127k_H 100,000 127,029 46
RandomBand_p3_b42_100k_127k_I 100,000 127,475 39
RandomBand_p3_b42_100k_127k_J 100,000 127,275 62
RandomBand_p5_b20_100k_102k_A 100,000 102,053 1,298
RandomBand_p5_b20_100k_102k_B 100,000 102,621 1,063
RandomBand_p5_b20_100k_102k_C 100,000 102,021 1,298
RandomBand_p5_b20_100k_102k_D 100,000 102,968 1,075
RandomBand_p5_b20_100k_102k_E 100,000 102,650 952
RandomBand_p5_b20_100k_102k_F 100,000 102,309 1,162
RandomBand_p5_b20_100k_102k_H 100,000 102,324 1,190
RandomBand_p5_b20_100k_102k_I 100,000 102,465 1,369
RandomBand_p5_b20_100k_102k_J 100,000 102,244 1,010
RandomBand_p5_b20_100k_103k_G 100,000 103,152 892
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics from SuiteSparse Matrix Collection [DH11] symmetrically permuted using the fill-reducing method ‘METIS_NodeND’ of [KK98]. The average wavefront size has been rounded down., label = table:hdagg-metis-graph-table, ] colspec = Q[l,m] | Q[r,m] | Q[r,m] | Q[r,m] , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Dimension #Non-zeros Avg. wavefront
af_0_k101_metis 503,625 9,027,150 610
af_shell10_metis 1,508,065 27,090,195 1,065
apache2_metis 715,176 2,766,523 47,678
audikw_1_metis 943,695 39,297,771 1,734
bmwcra_1_metis 148,770 5,396,386 473
bone010_metis 986,703 36,326,514 1,326
boneS10_metis 914,898 28,191,660 2,401
bundle_adj_metis 513,351 10,360,701 11,407
cant_metis 62,451 2,034,917 333
consph_metis 83,334 3,046,907 247
crankseg_2_metis 63,838 7,106,348 86
ecology2_metis 999,999 2,997,995 62,499
Emilia_923_metis 923,136 20,964,171 2,107
Fault_639_metis 638,802 14,626,683 1,458
Flan_1565_metis 1,564,794 59,485,419 2,569
G3_circuit_metis 1,585,478 4,623,152 93,263
Geo_1438_metis 1,437,960 32,297,325 2,887
gyro_metis 17,361 519,260 88
hood_metis 220,542 5,494,489 984
Hook_1498_metis 1,498,023 31,207,734 4,059
inline_1_metis 503,712 18,660,027 1,549
ldoor_metis 952,203 23,737,339 4,858
m_t1_metis 97,578 4,925,574 268
msdoor_metis 415,863 10,328,399 1,856
nasasrb_metis 54,870 1,366,097 287
PFlow_742_metis 742,793 18,940,627 1,023
pwtk_metis 217,918 5,926,171 511
raefsky4_metis 19,779 674,195 111
ship_003_metis 121,728 4,103,881 494
shipsec8_metis 114,919 3,384,159 456
StocF-1465_metis 1,465,137 11,235,263 11,446
thermal2_metis 1,228,045 4,904,179 45,483
tmt_sym_metis 726,713 2,903,837 26,915
x104_metis 108,384 5,138,004 306